# Representing Molecules with Algebraic Data Types: Beyond SMILES and SELFIES

### v4

@@ 3.9 Limitations and Future Work
The statistical and computational efficiency of our reference implementation has not been optimised for. Our current prototype prioritises clarity over compactness. In practice, the ADT is intended as an in-memory semantic representation; serialisation can target standard compact encodings (SDF/SMILES/InChI) or a dedicated binary format. We therefore treat file-size considerations as an engineering concern rather than a representational limitation. Furthermore, by storing a molecule as function or properties as higher-order functions, it may be possible to significantly reduce the storage size of a molecule.
@@ 3.10 Future work
The ADT so far presented has the ability to be improved upon, extended, and used in empirical experiments to verify and test the representation’s utility, ease-of-use and verifiability across a number of domains in cheminformatics. Parsers from other formats e.g. SMILES strings to this ADT could also be used. Currently only parsers for .SDF files have been written.
@@ 3.10.1 Geometric Deep Learning on Molecular Representations
Future work could consist of using symmetry-aware with Geometric Deep Learning, to reduce estimation error (requiring less data to infer parameters). A group captures the set of symmetrical operations over the representation. For molecules the key properties are (i) permutations of atom IDs and (ii) rigid 3D motions (rotations/translations). Randomized SMILES already tries to mimic (i) by enumerating alternative traversals; GDL would make this automatic—any relabeling or rigid motion is handled identically (invariant for scalar properties, equivariant for vector fields). Practically, we can encode this contract in code by introducing a small Haskell Group typeclass, with instances for atom-index permutations and rigid motions; models then declare which actions they respect, and the type system enforces symmetry-safe use. This process of learning leverages algebraic structure in the domain, to reduce estimation, model and approximation error amongst other things.
3.10.2 Extending and Evaluating the ADT
The abstract data type (ADT) presented here can be extended and empirically evaluated to test its utility across diverse tasks in cheminformatics. Two complementary lines of work are: (i) improving interoperability through additional parsers, and (ii) augmenting the core representation to cover chemical phenomena not yet encoded.
@@ 3.10.2 Extending and Evaluating the ADT
Empirical evaluation.
With these extensions and additional parsers in place, the ADT can be stress-tested across representative cheminformatics benchmarks to quantify downstream impact and validate its practical utility.
@@ 3.10.7 Experimental validation of the ADT
Neural networks can be designed to be invariant to permutations of atomic indexes, ensuring that molecularly equivalent structures are treated identically regardless of labeling of indices or bonds. Similarly, molecular features such as the atomSymbol, are both translation and rotation invariant, while others may only be translation invariant. Incorporating these symmetries directly into the neural network architecture could significantly reduce the amount of training data required and improve the generalization capability of the models, but this would have to be demonstrated.
@@ 4 Conclusion
In this work, we introduced a molecular representation grounded in algebraic data types and provided a reference implementation to make the method concrete and reproducible. The central idea is to treat molecules as structured, typed values—rather than as strings—so that validation, transformation, and composition are defined directly over the molecular structure and many malformed manipulations can be rejected early.
@@ 5.1 Availability and requirements
Project name: MolADT-Bayes (reference implementation for this paper)
@@
Tested on macOS Tahoe 26.2 (Build 24F74) (GHC 9.6.5, Cabal 3.14.2.0, Stack 3.7.1, ghcup 0.1.50.2)
@@
GHC via Stack; build with stack build; run stack exec moladtbayes. A step-by-step README.md documents OS, dependencies, and exact commands.
@@
Use of AI-assisted editing
We used OpenAI’s GPT-5 Pro (ChatGPT) only for light editorial assistance (copy-editing, LaTeX hygiene, float placement). All domain content—mathematics, chemistry, Haskell code, experiments, and conclusions—was written by the authors; some code refactoring drew on OpenAI’s Codex. Any LLM suggestions were edited, verified, and cited. No proprietary or human-subject data were provided. The authors take full responsibility for the manuscript and codebase.
@@ 5.2 Competing interests
The authors declare no competing interests. The repository is already listed on an open source molecular modeling index on GitHub [93].
@@ 5.3 Funding
Oliver Goldstein was funded in part by an EPSRC DTP Scholarship and is also partly self-funded.
Sam March is self-funded.
@@ 5.4.1 Oliver Goldstein
Ideated the project and conceptualized the solution in the form of an Algebraic Data Type, implemented the reference prototype and the worked examples/experiments. Oliver drafted the initial article and finalized its writing.

### v5

@@ 3.8 Comparison to existing tools
3.8.1 Representation-aware benchmarking.
When evaluating molecular ML models, it is rarely sufficient to report only dataset, architecture, and metric. The representation defines additional experimental conditions that should be held fixed—or at least reported—because they change the learning problem. At minimum, benchmarks should specify:
  • Validity contract: what structures are considered valid and how validity is enforced (parser-only, constrained decoding, rejection sampling, or representation-level invariants);
  • Edit locality: what constitutes a “small change” (token edit, graph edit, or structured transformation) and how perturbations are generated;
  • Symmetry handling: whether the model is required to be invariant/equivariant to atom-index permutations and (when 3D is present) rigid motions, and how this is enforced or tested.
MolADT is intended as a substrate that makes these explicit. By providing a typed semantic core with deterministic validation and well-scoped transformations, MolADT enables benchmarking studies that can ask representation-level questions directly for example, how model validity rates, novelty/diversity measures, uncertainty calibration, or sample efficiency change when the edit operations are local structure-preserving transformations rather than token mutations.
@@ 3.9 Limitations and Future Work
The statistical and computational efficiency of our reference implementation is explicitly not the focus of our article. Our current prototype prioritises clarity and correctness, over compactness. In practice, the ADT is intended as an in-memory representation; serialisation can target standard compact encodings (SDF/SMILES/InChI) or a dedicated binary format. We therefore treat file-size considerations as an engineering concern rather than a representational limitation. Furthermore, by storing a molecule as intensionally as a function or properties as higher-order functions, rather than extensionally, it may be possible to significantly reduce the storage size of a molecule.
@@ 3.10 Future work
The ADT so far presented has the ability to be improved upon, extended, and used in empirical experiments to verify and test the representation’s utility, ease-of-use and verifiability across a number of domains in cheminformatics. Parsers from other formats e.g. SMILES strings to this ADT could also be used. Currently a parser for .SDF files has been implemented.
@@ 3.10.1 Geometric Deep Learning on Molecular Representations
Future work could consist of using symmetry-aware with Geometric Deep Learning, to reduce estimation error (requiring less data to infer parameters). A group captures the set of symmetrical operations over the representation. For molecules the key properties are (i) permutations of atom IDs and (ii) rigid 3D motions (rotations/translations). Randomized SMILES already tries to mimic (i) by enumerating alternative traversals; GDL would make this automatic any relabeling or rigid motion is handled identically (invariant for scalar properties, equivariant for vector fields). Practically, we can encode this contract in code by introducing a small Haskell Group typeclass, with instances for atom-index permutations and rigid motions; models then declare which actions they respect, and the type system enforces symmetry-safe use. This process of learning leverages algebraic structure in the domain, to reduce estimation, model and approximation error amongst other things.
3.10.2 Extending and Conceptual Evaluation of the ADT
The Algebraic Data Type presented here can be extended and empirically evaluated to test its utility across diverse tasks in cheminformatics. Two complementary lines of work are: (i) improving interoperability through additional parsers, and (ii) augmenting the core representation to cover chemical phenomena not yet encoded.
@@ 3.10.6 Exploring the ADT Further
Neural networks can be designed to be invariant to permutations of atomic indexes, ensuring that molecularly equivalent structures are treated identically regardless of labeling of indices or bonds. Similarly, molecular features such as the atomSymbol, are both translation and rotation invariant, while others may only be translation invariant. Incorporating these symmetries directly into the neural network architecture could significantly reduce the amount of training data required and improve the generalization capability of the models.
@@ 4 Conclusion
In this work, we introduced a molecular representation grounded in Algebraic Data Types and provided a reference implementation to make the method concrete and reproducible. The central idea is to treat molecules as structured, typed values rather than as strings so that validation, transformation, and composition are defined directly over the molecular structure and many malformed manipulations can be rejected early. The contribution is therefore not a claim of new state-of-the-art task performance, but a reference representation and implementation intended to enable more controlled, interpretable, and reproducible benchmarking studies in cheminformatics.
@@ 5.1 Availability and requirements
Project name: MolADT-Bayes (reference implementation)
@@
Tested on macOS Tahoe 26.2 (GHC 9.6.5, Cabal 3.14.2.0, Stack 3.7.1, ghcup 0.1.50.2)
@@
GHC via Stack; build with stack build; run stack exec moladtbayes. A step-by-step README.md on the GitHub documents OS, dependencies, and exact commands.
@@ 5.2 Competing interests
The authors declare no competing interests. The repository is listed on an Open Source Molecular Modeling Index on GitHub [98].
@@ 5.3 Funding
Oliver Goldstein was funded in part by an EPSRC DTP Scholarship and is also partly self-funded. Sam March is self-funded.
@@ 5.4.1 Oliver Goldstein
Ideated the project and conceptualized the solution in the form of an Algebraic Data Type, implemented the reference prototype and the worked examples/experiments, and drafted the initial article.
@@ 20 Coordinate (mkAngstrom x) (mkAngstrom y) (mkAngstrom z)
1module Ferrocene (ferrocenePretty) where
2
3import qualified Data.Map.Strict as M
4import qualified Data.Set as S
5
6import Chem.Dietz
7 ( AtomId(..)
8 , SystemId(..)
9 , NonNegative(..)
10 , mkEdge
11 , mkBondingSystem
12 )
13import Chem.Molecule
14 ( AtomicSymbol(..)
15 , Molecule(..)
16 , Atom(..)
17 , Coordinate(..)
18 , mkAngstrom
19 )
20import Constants (elementAttributes, elementShells)
21
22ferrocenePretty :: Molecule
23ferrocenePretty = Molecule
24 { atoms = atomTable
25 , localBonds = sigmaFramework
26 , systems =
27 [ (SystemId 1, cp1PiSystem)
28 , (SystemId 2, cp2PiSystem)
29 , (SystemId 3, feBackDonationSystem)
30 ]
31 }
32 where
33 -- Atom IDs
34 fe = AtomId 1
35 ring1C = AtomId <\$> [2..6]
36 ring2C = AtomId <\$> [7..11]
37 ring1H = AtomId <\$> [12..16]
38 ring2H = AtomId <\$> [17..21]
39
40 -- Shared element data
41 feAttributes = elementAttributes Fe
42 carbonAttributes = elementAttributes C
43 hydrogenAttributes = elementAttributes H
44
45 feShells = elementShells Fe
46 carbonShells = elementShells C
47 hydrogenShells = elementShells H
48
49 -- Helpers
50 coord (x,y,z) =
51 Coordinate (mkAngstrom x) (mkAngstrom y) (mkAngstrom z)
52
53 mkAtom aid attrs sh xyz = Atom
54 { atomID = aid
55 , attributes = attrs
56 , coordinate = coord xyz
57 , shells = sh
58 , formalCharge = 0
59 }
60
61 mkEdges = S.fromList . map (uncurry mkEdge)
62
63 ringPairs xs = zip xs (tail (cycle xs))
64
65 -- (Replace with PubChem SDF coords if desired.)
66 feCoord = (0.0000, 0.0000, 0.0000)
67
68 ring1CarbonCoords =
69 [ ( 1.1800, 0.0000, 1.6600)
70 , ( 0.3647, 1.1220, 1.6600)
71 , (-0.9547, 0.6935, 1.6600)
72 , (-0.9547, -0.6935, 1.6600)
73 , ( 0.3647, -1.1220, 1.6600)
74 ]
75
76 ring2CarbonCoords =
77 [ ( 0.9547, 0.6935, -1.6600)
78 , (-0.3647, 1.1220, -1.6600)
79 , (-1.1800, 0.0000, -1.6600)
80 , (-0.3647, -1.1220, -1.6600)
81 , ( 0.9547, -0.6935, -1.6600)
82 ]
83
84 ring1HydrogenCoords =
85 [ ( 2.2700, 0.0000, 1.6600)
86 , ( 0.7016, 2.1582, 1.6600)
87 , (-1.8364, 1.3338, 1.6600)
88 , (-1.8364, -1.3338, 1.6600)
89 , ( 0.7016, -2.1582, 1.6600)
90 ]
91
92 ring2HydrogenCoords =
93 [ ( 1.8364, 1.3338, -1.6600)
94 , (-0.7016, 2.1582, -1.6600)
95 , (-2.2700, 0.0000, -1.6600)
96 , (-0.7016, -2.1582, -1.6600)
97 , ( 1.8364, -1.3338, -1.6600)
98 ]
99
100 feAtom = mkAtom fe feAttributes feShells feCoord
101
102 ring1CarbonAtoms =
[…]
104 ring1C ring1CarbonCoords
105
106 ring2CarbonAtoms =
[…]
108 ring2C ring2CarbonCoords
109
110 ring1HydrogenAtoms =
[…]
112 ring1H ring1HydrogenCoords
113
114 ring2HydrogenAtoms =
[…]
116 ring2H ring2HydrogenCoords
117
118 allAtoms = feAtom : (ring1CarbonAtoms ++ ring2CarbonAtoms ++ ring1HydrogenAtoms ++ ring2HydrogenAtoms)
119
120 atomTable = M.fromList [(atomID a, a) | a <- allAtoms]
121
122 -- adjacency (localised bonds): C-C rings + C-H
123 ring1CCPairs = ringPairs ring1C
124 ring2CCPairs = ringPairs ring2C
125 ring1CHPairs = zip ring1C ring1H
126 ring2CHPairs = zip ring2C ring2H
127
128 sigmaFramework =
129 mkEdges (ring1CCPairs ++ ring2CCPairs ++ ring1CHPairs ++ ring2CHPairs)
130
131 -- Dietz-style bonding systems (electron pools)
132 feToRing1 = [(fe, c) | c <- ring1C]
133 feToRing2 = [(fe, c) | c <- ring2C]
134 feToAll = feToRing1 ++ feToRing2
135
136 cp1Edges = mkEdges (feToRing1 ++ ring1CCPairs)
137 cp2Edges = mkEdges (feToRing2 ++ ring2CCPairs)
138 feBackEdges = mkEdges feToAll
139
140 cp1PiSystem =
141 mkBondingSystem (NonNegative 6) cp1Edges (Just "cp1_pi")
142
143 cp2PiSystem =
144 mkBondingSystem (NonNegative 6) cp2Edges (Just "cp2_pi")
145
146 feBackDonationSystem =
147 mkBondingSystem (NonNegative 6) feBackEdges (Just "fe_backdonation")
Listing A.2: Diborane (B2H6) with two 3c–2e bridges as Dietz pools.
1module Diborane (diboranePretty) where
2
3import qualified Data.Map.Strict as M
4import qualified Data.Set as S
5
6import Chem.Dietz
7 ( AtomId(..)
8 , SystemId(..)
9 , NonNegative(..)
10 , mkEdge
11 , mkBondingSystem
12 )
13import Chem.Molecule
14 ( AtomicSymbol(..)
15 , Molecule(..)
16 , Atom(..)
17 , Coordinate(..)
18 , mkAngstrom
19 )
20import Constants (elementAttributes, elementShells)
21
22diboranePretty :: Molecule
23diboranePretty = Molecule
24 { atoms = atomTable
25 , localBonds = sigmaFramework
26 , systems =
27 [ (SystemId 1, bridgeH3System)
28 , (SystemId 2, bridgeH4System)
29 ]
30 }
31 where
32 -- Atom IDs
33 b1 = AtomId 1
34 b2 = AtomId 2
35
36 h3 = AtomId 3 -- bridge
37 h4 = AtomId 4 -- bridge
38
39 h5 = AtomId 5 -- terminal on b1
40 h6 = AtomId 6 -- terminal on b1
41 h7 = AtomId 7 -- terminal on b2
42 h8 = AtomId 8 -- terminal on b2
43
44 -- Shared element data
45 boronAttributes = elementAttributes B
46 hydrogenAttributes = elementAttributes H
47
48 boronShells = elementShells B
49 hydrogenShells = elementShells H
50
51 -- Helpers
52 coord (x,y,z) =
53 Coordinate (mkAngstrom x) (mkAngstrom y) (mkAngstrom z)
54
55 mkAtom aid attrs sh xyz = Atom
56 { atomID = aid
57 , attributes = attrs
58 , coordinate = coord xyz
59 , shells = sh
60 , formalCharge = 0
61 }
62
63 mkEdges = S.fromList . map (uncurry mkEdge)
64
65 -- Idealised D2h-like geometry, chosen to make bridges explicit.
66 bCoords =
67 [ (-0.8850, 0.0000, 0.0000) -- B1
68 , ( 0.8850, 0.0000, 0.0000) -- B2
69 ]
70
71 hCoords =
72 [ ( 0.0000, 0.0000, 0.9928) -- H3 bridge
73 , ( 0.0000, 0.0000, -0.9928) -- H4 bridge
74 , (-0.8850, 1.1900, 0.0000) -- H5 terminal (B1)
75 , (-0.8850, -1.1900, 0.0000) -- H6 terminal (B1)
76 , ( 0.8850, 1.1900, 0.0000) -- H7 terminal (B2)
77 , ( 0.8850, -1.1900, 0.0000) -- H8 terminal (B2)
78 ]
79
80 boronAtoms =
[…]
82 [b1,b2]
83 bCoords
84
85 hydrogenAtoms =
[…]
87 [h3,h4,h5,h6,h7,h8]
88 hCoords
89
90 allAtoms = boronAtoms ++ hydrogenAtoms
91 atomTable = M.fromList [(atomID a, a) | a <- allAtoms]
92
93 -- adjacency: B-B and four terminal B-H bonds
94 sigmaFramework =
95 mkEdges [ (b1,b2)
96 , (b1,h5), (b1,h6)
97 , (b2,h7), (b2,h8)
98 ]
99
100 -- 3c-2e bridges as Dietz pools (2 electrons shared over two edges)
101 bridgeH3Edges = mkEdges [(b1,h3), (b2,h3)]
102 bridgeH4Edges = mkEdges [(b1,h4), (b2,h4)]
103
104 bridgeH3System =
105 mkBondingSystem (NonNegative 2) bridgeH3Edges (Just "bridge_h3_3c2e")
106
107 bridgeH4System =
108 mkBondingSystem (NonNegative 2) bridgeH4Edges (Just "bridge_h4_3c2e")
Listing A.3: Dietz-style bonding systems (paper): localized C–H and C–C bonds in localBonds ($\sigma$ adjacency), 6e pool over (Fe–C + ring C–C) for each Cp ring, 6e pool over all Fe–C edges